Not logged in
Thu Jun 16 18:08:57 2016, Johannes, Update, PSL, Before the AOM installation
Fri Jun 17 22:36:11 2016, Johannes, Update, PSL, AOM installation
Mon Jun 20 08:26:56 2016, Steve, Update, PSL, AOM pictures
Fri Jun 17 22:36:11 2016
In reply to:
Reply to this:
Subham and I have placed the AOM back into the setup right in front of the PMC.
The HEPA filters were turned off for some reason. They were turned back on, running at 100% while the enclosure was open.
Before the installation, after initial realignment, the PMC TRANSPD read out 748 mV.
The laser injection current was dialed down to 0.8 A (just above the threshold, judging by PMC cameras.
AOM was attached to the new mount while staying connected to its driver. Put in place, a clamp prevents the cable from moving anywhere near the main beam.
Aligned AOM to beam, centering the beam (by eye) on front and back apertures.
We then applied an offset to the AOM driver input, eventually increasing it to 0.5 V. A secondary beam became clearly visible
the primary beam.
In order to place the razor blade dump (stemming from a box, labeled "cleaned for atm use") before the PMC, where the beam separation was about 3 mm, to make sure we can hit the edged area, we had to place the dump at an angle, facing up.
Keeping the 0.5V offset on the driver input, with the lights off, we increased the laser diode current in steps of ~200 mA to its original value of 2.1A, while checking for any IR light scattered from the beam dump. Not a trace.
At original current setting, we realigned the beam into the PMC, and obtained 743 mV on the TRANSPD in the locked state.
Closed off PSL table, dialed HEPAs down to 50%